OpenStudy (anonymous):

Solve: negative 1 over 2 x + 2 = −x + 7 10 5 over 2 2 over 5 20

OpenStudy (anonymous):

Is this the original problem? \[\frac{-1 }{ 2x+2 } = -x+7\]

OpenStudy (anonymous):

Multiply both sides by (2x+2) to get rid of the fractions \[\frac{ -1 }{ 2x+2 }(2x+2)=(-x+7)(2x+2)\]

OpenStudy (anonymous):

\[Distribute (-2x^2-2x+7x+14)=-2x^2+5x+14\] Add one to both sides \[-1+(1)=(-2x^2+5x+14+(1))\] \[-2x^2+5x+15=0 \] Now factor and solve

OpenStudy (anonymous):

its it ten

OpenStudy (anonymous):

no, can you factor?

OpenStudy (anonymous):

I can help

OpenStudy (anonymous):

ok it looks like its 20

OpenStudy (anonymous):

My bad it was \[-2x^2+12x+14=-1 \] \[-2x^2+12x+15=0\]

OpenStudy (anonymous):

it was 10

OpenStudy (anonymous):

what was the question? the one iw rote up there?

OpenStudy (anonymous):

what

OpenStudy (anonymous):

what was the original thing casue it look confusing